In a displacement (m) vs time (s) graph, what is the velocity of an object, whose motion is shown with the equation y = (1/2)x?

A. - 2 m/s
B. 4 m/s
C. + 2 m/s
D. + 0.5 m/s
E. - 0.5 m/s

A bottle rocket is launched into the air with an initial velocity of 150 m/s. How long will it be in the air?

A. 5.25 seconds
B. 15 seconds
C. 30 seconds
D. 7.5 seconds

A rhino charges full speed at a car with an initial velocity of 15 m/s. When the rhino collides with the car, it crumples in by 1 m before the rhino comes to a complete stop. What acceleration did the rhino feel as it came to a stop?

A. - 112.5 m/s^2
B. - 7.5 m/s^2
C. - 30 m/s^2
D. + 112.5 m/s^2
E. + 30 m/s^2
F. + 7.5 m/s^2
1
Expert's answer
2020-10-13T09:36:30-0400

1) The velocity of the object, whose motion is shown with the equation y = (1/2)x, according to the similarity with d=vt, is +0.5 m/s.


2) A bottle rocket launched into the air with an initial velocity of 150 m/s will fall to the ground after


"T=2t=2\cdot v/g=2\cdot150/9.8=31\text{ s}."

3) The acceleration the rhino feels as it came to a stop can be found with a very handy equation:


"a=\frac{v_\text{final}-v_\text{iniutial}}{2\cdot\text{distance}},\\\space\\
a=\frac{0-15}{2\cdot1}=-7.5\text{ m/s}^2."